OpenStudy (anonymous):

csc(x) = −√2 find 4 x intercepts

OpenStudy (anonymous):

\[\sin \theta=\frac{ - 1}{ \sqrt{2} }\]

OpenStudy (anonymous):

so would that be -90 degrees and 270 degrees?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

is int 135 and 215

OpenStudy (anonymous):

both minus

OpenStudy (anonymous):

\[\theta=\sin ^{-1}\frac{ -1 }{ \sqrt{2} }+360k,k \in \mathbb{Z}\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

wait it cant be 135 because 135 is in the second quadrant

OpenStudy (anonymous):

225

OpenStudy (anonymous):

yes and 315

OpenStudy (anonymous):

both and negative would be 45 and 135

OpenStudy (anonymous):

positive yes